# One-Shot Manipulation Strategy Learning by Making Contact Analogies

### v1

@@ II Related Work
Our algorithm is inspired by contact-based modeling techniques used in robotic manipulation. Various approaches have been developed for planning manipulations in contact space, involving both rigid bodies and robotic hands [14, 15, 16, 17, 18, 19, 20, 21, 22]. Techniques such as CMGMP [23, 1, 2] and the work of [3] perform search over the exponential space of possible contact sequences to determine possible contact modes using three basic types: fixed, separating, and sliding contact, while [4] learns the types of the contact sequence from a single demonstration to guide the planning around them. By contrast, in this paper, we focus on leveraging global and local shape matching to generate analogical contact points using a single demonstration, which significantly improves the runtime efficiency of existing methods as well as the generalizations to unseen objects.
Recently, researchers have explored learning object affordances [24, 25, 26, 9, 11, 27, 28], usually from a large set of demonstrations. In contrast, this paper focuses on generating trajectories with novel objects from a single demonstration leveraging pretrained visual features and shape analysis algorithms, requiring no additional training data.
For many years, people have attempted to harness shape information to guide tool usage through mathematical analysis [29, 30, 31, 32, 33] and data-driven approaches [5, 34, 35, 36, 10, 37]. These methods often necessitate intricate human specifications or extensive in-domain datasets. In contrast, our approach integrates the off-the-shelf visual models, pretrained on general image datasets, with the generic geometrical property of curvature. This combination allows us to achieve effective and efficient tool-using guidance with minimal human intervention and without the need for large specialized datasets.
@@ III-A Problem Definition
Overall, the input to magic is an image $I_{T}$ of the reference object $T$, an image $I_{T^{\prime}}$ of the target object, and a point of interest $p_{T}$ on the reference image. Our goal is to find the corresponding point $p_{T^{\prime}}$ on $I_{T^{\prime}}$. When we have two objects interacting: the tool object $T$ being directly held by the robot (e.g., the hook) and another object $O$ in contact with $T$ (e.g., the target object we want to hook), our goal would be to find a contact point pair $p_{T^{\prime}}$ on $I_{T^{\prime}}$ and $p_{O^{\prime}}$ on $I_{O^{\prime}}$ given the reference $\left(T,O,p_{T},p_{O}\right)$ that maximizes a score function: $\textit{score}\left(X,X^{\prime}\right)$, where $X=\left(T,O,p_{T},p_{O}\right)$ is the reference contact and $X^{\prime}=\left(T^{\prime},O^{\prime},p_{T^{\prime}},p_{O^{\prime}}\right)$ is the target contact. In this paper, we employ a two-stage global-to-local matching process, therefore, the score function will be composed of two parts: a global matching score $s_{\textit{dino}}$, and a local matching score $s_{\textit{curv}}$:
$\textit{score}\left(X,X^{\prime}\right)=s_{\textit{dino}}(T,T^{\prime},p_{T},p_{T^{\prime}})+\lambda\cdot s_{\textit{curv}}(X,X^{\prime}),$
@@ III-C Local Contact Point Matching with Curvature Estimation
$s_{\textit{curv}}\left(\langle T,O,p_{T},p_{O}\rangle,\langle T^{\prime},O^{\prime},p_{T^{\prime}},p_{O^{\prime}}\rangle\right)=\left|\frac{r(p_{T})}{r(p_{O})}-\frac{r(p_{T^{\prime}})}{r(p_{O^{\prime}})}\right|.$
Figure 4: (a) The observing scale at which we perform estimation has effect on both the magnitude and the sign of he estimated curvature. (b) Filtering out the irrelevant points can give a more accurate curvature estimation. (c) We perform local search with in the observation region to find the contact point with the correct convexity.
@@ V Conclusion
Acknowledgements. We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; and from ARO grant W911NF-23-1-0034; from MIT Quest for Intelligence; from the MIT-IBM Watson AI Lab; from ONR Science of AI; and from Simons Center for the Social Brain. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.
@@
Figure 10: Example episodes of Hanging and Hooking in the real world.

### v2

@@ II Related Work
Contact-Based Modeling in Robotic Manipulation. Various approaches have been developed for planning manipulations in contact space, involving both rigid bodies and robotic hands [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Techniques such as CMGMP [25, 1, 2] and the work of [3] perform search over the exponential space of possible contact sequences to determine possible contact modes using three basic types: fixed, separating, and sliding contact, while [4] learns the types of the contact sequence from a single demonstration to guide the planning around them. By contrast, in this paper, we focus on leveraging global and local shape matching to generate analogical contact points using a single demonstration, which significantly improves the runtime efficiency of existing methods as well as the generalizations to unseen objects.
Affordance Learning and Keypoint Representations. Object affordance characterizes how an object can interact with others, often represented through keypoints. Traditionally, this concept has been learned from large dataset  [26, 27, 28, 9, 10, 11, 29]. In contrast, our approach generates trajectories for novel objects from a single demonstration without requiring additional training data. More recent works [30, 31, 32, 33, 34] have explored one-shot or zero-shot affordance using vision and language foundation models; however, these methods operate primarily within the semantic domain and may not generalize well to creative, tool-using scenarios. Our work distinguishes itself by proposing precise, physically plausible contact points derived from pretrained visual features and advanced shape analysis algorithms.
Shape-Guided Tool Using. For many years, people have attempted to harness shape information to guide tool usage through mathematical analysis [35, 36, 37, 38, 39] and data-driven approaches [5, 40, 41, 42, 10, 43]. These methods often necessitate intricate human specifications or extensive in-domain datasets. In contrast, our approach integrates the off-the-shelf visual models, pretrained on general image datasets, with the generic geometrical property of curvature. This combination allows us to achieve effective and efficient tool-using guidance with minimal human intervention and without the need for large specialized datasets.
@@ III-A Problem Definition
Overall, the input to magic is an image $I_{T}$ of the reference object $T$, an image $I_{T^{\prime}}$ of the target object, and a point of interest $p_{T}$ on the reference image. Our goal is to find the corresponding point $p_{T^{\prime}}$ on $I_{T^{\prime}}$. When we have two objects interacting: the tool object $T$ being directly held by the robot (e.g., the hook) and another object $O$ in contact with $T$ (e.g., the target object we want to hook), our goal would be to find a contact point pair $p_{T^{\prime}}$ on $I_{T^{\prime}}$ and $p_{O^{\prime}}$ on $I_{O^{\prime}}$ given the reference $\left(I_{T},I_{O},p_{T},p_{O}\right)$ that maximizes a score function: $\textit{score}\left(X,X^{\prime}\right)$, where $X=\left(I_{T},I_{O},p_{T},p_{O}\right)$ is the reference contact and $X^{\prime}=\left(I_{T^{\prime}},I_{O^{\prime}},p_{T^{\prime}},p_{O^{\prime}}\right)$ is the target contact. In this paper, we employ a two-stage global-to-local matching process, therefore, the score function will be composed of two parts: a global matching score $s_{\textit{dino}}$, and a local matching score $s_{\textit{curv}}$:
$\textit{score}\left(X,X^{\prime}\right)=s_{\textit{dino}}(I_{T},I_{T^{\prime}},p_{T},p_{T^{\prime}})+\lambda\cdot s_{\textit{curv}}(X,X^{\prime}),$
@@ III-C Local Contact Point Matching with Curvature Estimation
$s_{\textit{curv}}\left(\langle I_{T},I_{O},p_{T},p_{O}\rangle,\langle I_{T^{\prime}},I_{O^{\prime}},p_{T^{\prime}},p_{O^{\prime}}\rangle\right)=\left|\frac{r(p_{T})}{r(p_{O})}-\frac{r(p_{T^{\prime}})}{r(p_{O^{\prime}})}\right|.$
Figure 4: (a) The observing scale at which we perform estimation has effect on both the magnitude and the sign of the estimated curvature. (b) Filtering out the irrelevant points can give a more accurate curvature estimation. (c) We perform local search within the observation region to find the contact point with the correct convexity.
@@ V Conclusion
Acknowledgements. We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; and from ARO grant W911NF-23-1-0034; from MIT Quest for Intelligence; from the MIT-IBM Watson AI Lab; from the Boston Dynamics AI Institute; from ONR Science of AI; and from Simons Center for the Social Brain. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.
Figure 10: Example episodes of Hanging and Hooking in the real world.